Prev Next

SW&IT Meeting Minutes 2020-02-03

Attendees: Carl, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20200127
LR20200127 - Bec, Matthieu 40.0
LR20200127 - Holmberg, Carl 32.0
LR20200127 - Edmeades, David 40.0
LR20200127 - Sola, Igor 0.0
LR20200127 - Kubanek, Petr 0.0
LR20200127 - Hooper, Stephen 80.0
LR20200127 - Zhang, Yang 0.0
LR20200127 - Zhang, Xianyu 40.0

News from management

Previous Action Items

Owner Action Due Date Completion Date
DE Summit Veam Backup (soul-telemetry, etc) + bring back VM2 until PEPSI/VM arrives 2020-01-20  

Open topics

  • Remote Observing 'working team'
  • TMS development
  • rm507-x rollout E-time

Team reports

PK

  • AGw[34]
    • clioac improvements, documentation
  • 2020B
    • hope for today deployment, not testing until tomorrow (storm)

MB

  • Last week
    • OC, WPC, Management
    • board memo for new instruments s/w
    • DMS websocket/mongodb service
    • investigated few ITs (allsky @1sec, dimm no seeing values, lbc bluetech)
    • submitted s/w post 2
    • GMT review Monday
  • This week
    • DMS continued
    • sciarc webinar
    • hydrogen AHS
    • TCS ITs review with PK and YZ
    • retailor s/w post 1
    • Remote Observing kick-off
    • TMS kick-off

DE

  • PEPSI VM hosts setup
    • OVIRT install
    • Sample VM setup this week
  • Archive webinar
    • I have the power!
    • Test was partially successful this morning
    • Will follow up with Cristina
  • Old MBP battery diag and removal
    • Replacement this week
  • New safety training
  • KVM up to the mountain
    • Temporary, but how temporary? Do we need more dongles?
  • Peter Gray Autocad install
  • Veeam update quote
    • $2k for adding the mountain VM host
    • $5k if we reactivate the existing install for upgrades+add mountain host
  • Open conversation with Alexander about laptop

XZ

  • Last week
    • Power Backplane dicussion/review
    • config ao-analysis machine
    • organise AO telemetry files
  • This week
    • Power Backplane meeing with Robtero
    • attend SOUL commission?

SH

  • LBC issues
  • Telemetry positional modification
  • SOUL AGW, SOUL
  • info.lbto.org changes

next:

  • DNS Tucson

CH

Last week
  • AGW: Assisted PK in diagnosing AGW[3,4] UMAC data discontinuity during telemetry collection.
  • AGW: Assisted John Morris on getting a Remote Desktop client running on MacOS to access azcamserver.
  • AHS: Completed a beta laser power check script, need to add tms:state:laser EPICS channel to move data.
  • AHS: Queried LBTI on proper INDI signals and API to query data needed to add an alert.
  • TCS: Shadowed PK for TCS 2020B code changes.

This week
  • In late on Friday, doctor appointment.
  • AGW: Write IT to fix AGW firmware address conflict.
  • AHS: Deploy added HBS alert TCS changes.
  • ARGOS: work on IT6224, several AOS data items missing when ARGOS is operating.
  • Safety training.

IS

Last week:
  • Time Accounting Tool
    • Make report page be accessible via URL/REPL
    • Adding navigation to the report page
  • OT
    • Working with Andrew on OT bug fixes
This week:
  • Time Accounting Tool
    • Export to excel
    • Look into django admin login capabilities for the tool
  • OT
    • Finish bug fixes and release build.
Topic revision: r9 - 03 Feb 2020, DavidEdmeades
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback